A random variable X is given by its distribution table:

X -10 5 10 20

P .15 .4 .35 .1

Find its variance and standard deviation.
Answer #1

Solution:

x P(x) x * P(x) x2 * P(x)
-10 0.15 -1.5 15
5 0.4 2 10
10 0.35 3.5 35
20 0.1 2 40
Sum 1 6 100

Mean = X * P(X)

= 6

Standard deviation =

=X 2 * P(X) - 2

=  100 - (6)2 = 64

Standard deviation = = 8